# G2SF-MIAD: Geometry-Guided Score Fusion for Multimodal Industrial Anomaly Detection

### v1

@@ Abstract
Industrial quality inspection plays a critical role in modern manufacturing by identifying defective products during production. While single-modality approaches using either 3D point clouds or 2D RGB images suffer from information incompleteness, multimodal anomaly detection offers promise through the complementary fusion of crossmodal data. However, existing methods face challenges in effectively integrating unimodal results and improving discriminative power. To address these limitations, we first reinterpret memory bank-based anomaly scores in single modalities as isotropic Euclidean distances in local feature spaces. Dynamically evolving from Eulidean metrics, we propose a novel Geometry-Guided Score Fusion (G2SF) framework that progressively learns an anisotropic local distance metric as a unified score for the fusion task. Through a geometric encoding operator, a novel Local Scale Prediction Network (LSPN) is proposed to predict direction-aware scaling factors that characterize first-order local feature distributions, thereby enhancing discrimination between normal and anomalous patterns. Additionally, we develop specialized loss functions and score aggregation strategy from geometric priors to ensure both metric generalization and efficacy. Comprehensive evaluations on the MVTec-3D AD dataset demonstrate the state-of-the-art detection performance of our method with low positive rate and better recall, which is essential in industrial application, and detailed ablation analysis validates each component’s contribution. (Code will be released upon acceptance).
I Introduction
@@
In multimodal scenarios, a challenge centers on developing effective fusion strategies to integrate detection results from individual modalities. Despite efforts to cross-modal relationships, existing methods typically employ simplistic strategies such as direct addition [3] or maximum value selection [8] of anomaly scores across modalities, while [7] utilizes a One-Class Support Vector Machine (OCSVM) for score combination. However, these methodologies face a critical limitation: unimodal anomaly scores frequently demonstrate insufficient discriminative capacity in practical applications. As visualized by Fig. 1, even normal regions in both point cloud and image data exhibit unexpectedly elevated anomaly scores. These unreliable unimodal assessments consequently compromise the fusion performance. To obtain more discriminative anomaly scores, feature adaptation [9, 6] can be applied that adapts original features into more compact representations beforehand, followed by standard methods, e.g., PatchCore [5], to calculate final anomaly scores. One limitation is that critical details of original features essential for anomaly detection may be over-confidently compromised during first-step adaptation.
@@
Despite the enhanced discriminative power, our G2SF offers two other advantages: (i) Instead of learning a new metric or anomaly detector from scratch, which are often risky in high-dimensional feature spaces [10, 11], perhaps increasing with multiple modalities involved, G2SF explicitly quantifies how unimodal anomaly scores and local geometric structures in feature spaces contribute to the final metric. Consequently, it enables a progressive evolution from the Euclidean metric to an anisotropic metric, mitigating overfitting and ensuring robust metric learning (see Section III-D). (ii) As a one-stage anomaly detector, G2SF maintains complete modality-specific information throughout the detection process (details are discussed in Section III-C), in contrast to existing feature adaptation frameworks [9, 6] aforementioned above.
@@
  • We propose a novel metric deformation paradigm, evolving from isotropic Euclidean metric to an anisotripic one. Specifically, throughout geometric encoding, we design a new LSPN to predict direction-aware scaling factors, subsequently formulating the final metric.
  • We propose critical loss functions to maintain specific geometric properties (e.g., metric consistency and cross-modal correspondence) during metric learning and a novel geometry-informed anomaly scoring strategy, enabling precise multimodal anomaly detection.
  • Our method achieves the best performance on the MVTec 3D-AD dataset, with particularly notable superiority at the AUPRO@1% metric, which corresponds to a very low false positive rate and better recall. This is highly valuable in real industrial applications.
II Literature Review
II-A 2D Image Anomaly Detection
Unsupervised anomaly detection methodologies can be broadly categorized into two paradigms: reconstruction-based approaches and feature-based techniques. Reconstruction-based methods typically employ generative architectures, e.g., transformers [12] and diffusion models [13], to establish pixel-level reconstruction fidelity for normal images. During the inference phase, anomaly detection is performed by calculating per-pixel discrepancies between input images and their reconstructed counterparts.
@@ II Literature Review
II-B Multimodal 2D/3D Anomaly Detection
The pioneering MVTec-3D AD dataset [2] benchmarks this task with various generative networks. For real-time requirement, Easy-Net [8] designs a multi-scale multimodal autoencoder. Feature-based approaches also dominate in this field. CFM [18] constructs explicit one-to-one feature mappings to model the cross-modal relationship. For score fusion from different modalities, BTF [3] concatenates 2D/3D features with PatchCore [5], equivalent to linear score fusion. Based on the signed distance function (SDF) for 3D patch representation, Shape-Guided [19] establishes SDF-guided memory banks for fine-grained detection on image modality but relies on simplistic maximum value between image and SDF scores. M3DM [7] learns fused features by considering cross-modal interactions yet employs one-class SVM for data-driven score fusion of the original and fused features. However, the inherent limitations in the discriminative power of unimodal anomaly scores remain unaddressed. Toward this deficiency, LSFA [6] proposes a cross-modal feature adaptation paradigm that learns task-oriented features but risks overlooking the critical details of original features through adaptation.
@@ III-A Feature Extraction
III-B Prototype Retrieval from Memory Bank
Our G2SF builds on the memory bank-based methods for unimodal anomaly detection. For each modality $m\in{P,R}$, we construct a memory bank $\mathcal{M}^{m}$ that collects prototypical features from the entire training dataset following PatchCore [5]. Subsequently, for a feature $\mathbf{f}^{m}_{i}$, we can retrieve its $12k+$ 1 nearest neighbors $\{\mathbf{m}^{m}_{i,j}\}_{j=0}^{2k}$ from $\mathcal{M}^{m}$. Critically, the Euclidean distance between $\mathbf{f}^{m}_{i}$ and the nearest $\mathbf{m}^{m}_{i,0}$ is typically defined as the anomaly score $s^{m}_{i,0}$ in these memory bank-based methods [7].
III-C Geometric Feature Encoding
The main issue that limits the discriminative capability of $s^{m}_{i,0}$ in Section III-B is its isotropic nature, which ignores other critical characteristics valuable for anomaly detection. Nevertheless, local directional consistency can act as an auxiliary property to quantitatively evaluate the normality of target samples: Features aligned with the principal directions of normal samples within the local feature spaces are expected to demonstrate a higher degree of normality.
@@ III-C Geometric Feature Encoding
III-D Anisotropic Local Distance Metric
@@
The architecture of $\Phi(\cdot)$ is illustrated in Fig. 4, which consists of two parallel branches that independently concatenate and encode the prototypes $(\mathbf{m}^{P}_{i,j},\mathbf{m}^{R}_{i,j})$ and directions $(\mathbf{d}^{P}_{i,j},\mathbf{d}^{R}_{i,j})$. $w^{m}_{i,j}$ is outputed through an activation $\exp({\tanh{(\cdot)}})$ at the final layer, ensuring $w^{m}_{i,j}\in[e^{-1},e^{1}]$. This constrains the scaling factors to be symmetric around the Euclidean baseline ($w^{m}_{i,j}=1$). To enhance the discriminative power of $l(\cdot)$, we aim to reduce scales ($w^{m}_{i,j}<1$) for normal samples, while increasing scales ($w^{m}_{i,j}>1$) for anomalous samples.
III-E Loss Functions
Due to the lack of anomalous samples, the predicted $w^{m}_{i,j}$ will converge to $e^{-1}$ during training, yielding trivial direction-invariant $l(\cdot)$. We address this through synthetic anomaly injection in Section III-G. We denote the label $y_{i}\in\{0,1\}$ for $(\mathbf{f}^{P}_{i},\mathbf{f}^{R}_{i})$, where $y_{i}=1$ denotes synthetic anomalies. Additionally, we denote $(\cdot)_{+}$ as $\max(\cdot,0)$.
III-E1 Discrimination Loss
This loss aims to promote the discrepancy of learned metric values $l(\cdot)$ between normal ($y_{i}=0$) and anomalous ($y_{i}=1$) samples. Moreover, we prioritize $l_{i,0}$ to inherit its discriminative capability with respect to the nearest prototype $\bm{m}^{m}_{i,0}$.
Separation Loss: In line with the semi-supervised anomaly detection framework proposed in [25], the separation loss $\mathcal{L}_{sep}$ is designed to enforce the minimization of metric values for normal samples while simultaneously maximizing them for anomalous instances, formally expressed as:
@@ III-E1 Discrimination Loss
where the threshold $m_{0}$ prevents excessively large anomaly scores by setting $m_{0}=2\max_{i,m}{s^{m}_{i,0}}$ among all training samples, ensuring numerical stability.
Margin Loss: To address the boundary ambiguity issue not explicitly considered by $\mathcal{L}_{sep}$, the margin loss $\mathcal{L}_{mar}$ mitigates the distribution overlap between normal and anomalous samples, adapted from [26]:
@@
where $m_{l}=\min_{i|y_{i}=1}l_{i,0}$ is the lower bound of $l_{i,0}$ for anomalous samples and $m_{u}=\max_{i|y_{i}=0}l_{i,0}$ is the upper bound for normal samples. $m_{l}$ and $m_{u}$ are computed in a batch-wise.
III-E2 Geometry-Preserved Loss
Since learning a robust and generalizable $l(\cdot)$ in the high-dimensional space for anomaly detection is challenging, despite the architecture design in Section III-D, we aim to exploit the desired geometric properties of $l(\cdot)$ that benefit our task.
Figure 5: (a) Consistency loss $\mathcal{L}_{cns}$, where Proximal consistency enforces an upper bound on $l_{i,j}|^{k}_{j=1}$ to ensure metric continuity, and Distal separation imposes a lower bound on $l_{i,j}|^{2k}_{j=k+1}$ to avoid incredible values. (b) Cross-modal alignment loss $\mathcal{L}_{cma}$, where we disentangle cross-modal correspondences between prototypes and directions from a global-to-local granularity.
Consistency Loss: Eqs. (4) and (5) concern on the discriminative power of $l_{i,0}$, critical behaviors of $l_{i,j}$ across different local spaces around $\mathbf{m}^{m}_{i,j}$ remain unexploited. Incorporating $l_{i,j}$ serves two critical purposes: (i) Spatial regularization that stabilizes metric learning. (ii) Data augmentation that effectively increases sample diversity within each local space through multi-prototype observations. Specifically, we promote the following two properties of geometric consistency, as illustrated in Fig. 5:
@@ III-E2 Geometry-Preserved Loss
With the above purposes, $\mathcal{L}_{cns}$ is formulated as:
@@
We set $\eta_{i,j}=\eta_{0}\cdot(\sum_{m}s^{m}_{i,j})/(\sum_{m}s^{m}_{i,0})$ proportional to the original Euclidean distances.
@@
$\mathcal{L}_{sc}$ asymmetrically compresses scaling factors of normal samples while anchoring synthetic anomalies near the upper limit ($e^{1}$). This design is motivated by the observation that synthetic anomalies exhibit more pronounced deviations than real-world anomalies. Consequently, their penalization in $\mathcal{L}_{sc}$ is more aggressive.
Cross-modal Alignment Loss: So far, the prototypes (directions) of two modalities have been naively concatenated as input to LSPN in Section III-D, not considering cross-modal geometric alignments. To address it, we propose a simple yet effective strategy by disentangling cross-modal correspondences between prototypes and directions.
Specifically, by permuting indices $\{i\}\rightarrow\{\tilde{i}\}$, we can construct two kinds of negative samples, as illustrated in Fig. 5:
@@
The latter focuses on capturing discordant directions while maintaining prototype consistency at a local structural granularity; in contrast, the former prioritizes holistic prototype coherence through a global structural perspective.
During training, we construct the two kinds of negative samples with equal probability. Denote $\tilde{w}^{m}_{i,0}$ ($m\in\{P,R\}$) as the scaling factors corresponding to these negative samples, we define the cross-modal alignment loss as:
@@ III-E3 Overall Loss
III-F Geometry-Informed Anomaly Scoring
@@
This definition stems from a geometric prior. Specifically, memory bank-based methods calculate anomaly scores using Euclidean distances to their memory banks (Section III-B), whereas our formulation in Eq. (11) reformulates this as distance measurement to the normal data manifold through the anisotropic metrics $l(\cdot)$. We will compare Eq. (11) with several alternatives in Section IV-C.
@@ III-G Anomaly Synthesis
Ours 0.982 0.976 0.982 0.979 0.971 0.976 0.982 0.983 0.978 0.981 0.979
TABLE I: I-AUROC and AUPRO@30% results on MVTec-3D AD. Best results in bold, runner-ups underlined.
IV Experiment
@@
TABLE II: Mean metrics for all categories of MVTec-3D AD.
IV-A Experimental Details
Dataset: Experiments are conducted on the MVTec-3D AD dataset [2] containing 10 industrial object categories. Training data comprises 210–361 samples per class (total 2,656), while testing data contains 69–132 samples per class (total 249). Each sample provides aligned RGB images and 3D point clouds, enabling pixel-wise correspondence between color and spatial coordinates.
Evaluation Metrics: We assess the image-level anomaly detection performance by the Area Under the Receiver Operator Curve (I-AUROC). For evaluating the pixel-wise anomaly segmentation, we adopt the pixel-level AUROC (P-AUROC) and the Area Under the Per-Region Overlap (AUPRO) at 30% integration limits (denoted as AUPRO@30%), the same as existing methods [18, 3, 19, 7]. To better demonstrate the advantages of our method in practical industrial scenarios, particularly its low false positive rate and better recall, we have additionally introduced the AUPRO@1% metric used in recent studies [19, 18].
Implementation: We follow a similar procedure of M3DM [7] for feature extraction and preprocessing of the dataset. For the consistency loss $\mathcal{L}_{cns}$, we set $k=5$ and $\eta_{0}=1.2$. The tuning parameters associated with each loss are set as $\alpha=10,\beta=60,\gamma=8,\mu=20$. We train our G2SF with the Adam optimizer [30] with 80 epochs and batch size 8192. More details are provided in the supplementary material.
Figure 7: Pixel-wise anomaly score distributions in Bagel, Cable gland, Foam, and Rope categories, where the horizontal axis is normalized anomaly score (0-1) and the vertical axis is the probability density. Green/Brown distributions represent normal/anomalous regions respectively. Compared with CFM [18] (top), our method (bottom) achieves significant distribution compression in normal regions while maintaining anomaly discrimination.

### v2

# G2SF: Geometry-Guided Score Fusion for Multimodal Industrial Anomaly Detection

@@ Abstract
Industrial quality inspection plays a critical role in modern manufacturing by identifying defective products during production. While single-modality approaches using either 3D point clouds or 2D RGB images suffer from information incompleteness, multimodal anomaly detection offers promise through the complementary fusion of crossmodal data. However, existing methods face challenges in effectively integrating unimodal results and improving discriminative power. To address these limitations, we first reinterpret memory bank-based anomaly scores in single modalities as isotropic Euclidean distances in local feature spaces. Dynamically evolving from Euclidean metrics, we propose a novel Geometry-Guided Score Fusion (G2SF) framework that progressively learns an anisotropic local distance metric as a unified score for the fusion task. Through a geometric encoding operator, a novel Local Scale Prediction Network (LSPN) is proposed to predict direction-aware scaling factors that characterize first-order local feature distributions, thereby enhancing discrimination between normal and anomalous patterns. Additionally, we develop specialized loss functions and score aggregation strategy from geometric priors to ensure both metric generalization and efficacy. Comprehensive evaluations on the MVTec-3D AD and Eyecandies datasets demonstrate the state-of-the-art detection performance of our method, and detailed ablation analysis validates each component’s contribution. Our code is available at https://github.com/ctaoaa/G2SF. Corresponding author
1 Introduction
@@
Despite efforts in multimodal methods to learn intermodal representations or mining crossmodal relationships [1, 11, 32], leveraging the valuable unimodal detection results complementary to the above intermodal information remains underexploited. Current strategies—such as score addition [19], max selection [10], or One-Class Support Vector Machine (OCSVM) [32]—face a critical limitation: unimodal anomaly scores frequently demonstrate insufficient discriminative capacity in practical applications. As visualized by Fig. 1, even normal regions in both point cloud and image data exhibit unexpectedly elevated anomaly scores. These unreliable unimodal assessments consequently compromise the fusion performance. To obtain more discriminative anomaly scores, feature adaptation [26, 30] can be applied that adapts original features into more compact representations beforehand, followed by standard methods, e.g., PatchCore [27], to calculate final anomaly scores. One limitation is that critical details of original features essential for anomaly detection may be over-confidently compromised during first-step adaptation.
@@
Despite the enhanced discriminative power, our G2SF offers two other advantages: (i) Instead of learning a new metric or anomaly detector from scratch, which are often risky in high-dimensional feature spaces [15, 14], perhaps increasing with multiple modalities involved, G2SF explicitly quantifies how unimodal anomaly scores and local geometric structures in feature spaces contribute to the final metric. Consequently, it enables a progressive evolution from the Euclidean metric to an anisotropic metric, mitigating overfitting and ensuring robust metric learning (see Section 3.4). (ii) As a one-stage anomaly detector, G2SF maintains complete modality-specific information throughout the detection process (details are discussed in Section 3.3), in contrast to existing feature adaptation frameworks [26, 30] aforementioned above.
@@
  • We propose a novel metric deformation paradigm, evolving from isotropic Euclidean metric to an anisotropic one. Specifically, throughout geometric encoding, we design a new LSPN to predict direction-aware scaling factors, subsequently formulating the final metric.
  • We propose critical loss functions to maintain specific geometric properties (e.g., metric consistency and crossmodal correspondence) during metric learning and a novel geometry-informed anomaly scoring strategy, enabling precise multimodal anomaly detection.
  • Our method achieves state-of-the-art performance on the MVTec 3D-AD and Eyecandies datasets, with particularly notable superiority in the AUPRO@1% metric.
2 Literature Review
2.1 2D Image Anomaly Detection
Unsupervised anomaly detection methodologies can be broadly categorized into two paradigms: reconstruction-based approaches and feature-based techniques. Reconstruction-based methods typically employ generative architectures, e.g., transformers [25] and diffusion models [33], to establish pixel-level reconstruction fidelity for normal images. For inference, anomaly detection is performed by calculating per-pixel discrepancies between input images and their reconstructed counterparts.
@@ 2 Literature Review
2.2 Multimodal 2D/3D Anomaly Detection
The pioneering MVTec-3D AD dataset [4] benchmarks this task with various generative networks. For real-time requirements, Easy-Net [9] designs a multi-scale multimodal autoencoder. Feature-based methods prevail in this domain. CFM [11] establishes explicit one-to-one feature mappings to model crossmodal relationships, while 2M3DF [1] extends this by learning intermodal features via multi-view RGB image generation. Similarly, M3DM [32] incorporates mutual information for feature fusion. On the other hand, unimodal anomaly scores-containing essential modality-specific information-also demand dedicated fusion strategies. BTF [19] concatenates 2D/3D features with PatchCore [27], equivalent to linear score fusion. Based on the signed distance function (SDF) for 3D patch representation, Shape-Guided [10] establishes SDF-guided memory banks for fine-grained detection on image modality but relies on simplistic maximum value between image and SDF scores. M3DM [32] employs one-class SVM for data-driven score fusion of the original and fused features. However, the inherent limitations in the discriminative power of unimodal anomaly scores remain unaddressed. Toward this deficiency, LSFA [30] proposes a crossmodal feature adaptation paradigm that learns task-oriented features but risks overlooking the critical details of original features through adaptation. Finally, at data-level, DAUP [23] attributes the issue of indiscriminative point cloud scores to the irregular point density and tackles it by resampling point clouds.
@@ 3.1 Feature Extraction
3.2 Prototype Retrieval from Memory Bank
[…]
3.3 Geometric Feature Encoding
The main issue that limits the discriminative capability of $s^{m}_{i,0}$ in Section 3.2 is its isotropic nature, which ignores other critical characteristics valuable for anomaly detection. Nevertheless, local directional consistency can act as an auxiliary property to quantitatively evaluate the normality of target samples: Features aligned with the principal directions of normal samples within the local feature spaces are expected to demonstrate a higher degree of normality.
@@ 3.3 Geometric Feature Encoding
3.4 Anisotropic Local Distance Metric
@@
The architecture of $\Phi(\cdot)$ is illustrated in Fig. 4, which consists of two parallel branches that independently concatenate and encode the prototypes $(\mathbf{m}^{P}_{i,j},\mathbf{m}^{R}_{i,j})$ and directions $(\mathbf{d}^{P}_{i,j},\mathbf{d}^{R}_{i,j})$. $w^{m}_{i,j}$ is obtained through an activation $\exp({\tanh{(\cdot)}})$ at the final layer, ensuring $w^{m}_{i,j}\in[e^{-1},e^{1}]$. This constrains the scaling factors to be symmetric around the Euclidean baseline ($w^{m}_{i,j}=1$). To enhance the discriminative power of $l(\cdot)$, we aim to reduce scales ($w^{m}_{i,j}<1$) for normal samples, while increasing scales ($w^{m}_{i,j}>1$) for anomalous samples.
3.5 Loss Functions
To ensure stable training, we employ synthetic anomaly injection in Section 3.7. We denote the label $y_{i}\in\{0,1\}$ for $(\mathbf{f}^{P}_{i},\mathbf{f}^{R}_{i})$, where $y_{i}=1$ denotes synthetic anomalies. Additionally, we denote $(\cdot)_{+}$ as $\max(\cdot,0)$. We deploy the discrimination losses ($\mathcal{L}_{sep},\mathcal{L}_{mar}$) in Section 3.5.1 to ensure the learned $l(\cdot)$ capable of discriminating normal ($y_{i}=0$) and anomalous ($y_{i}=1$) samples. Furthermore, the geometry-preserved losses ($\mathcal{L}_{cns},\mathcal{L}_{sc},\mathcal{L}_{cma}$) are proposed in Section 3.5.2 to ensure effective metric learning by incorporating desired geometric properties related to metric consistency, local scaling factors, and crossmodal alignments.
3.5.1 Discrimination Loss
We prioritize $l_{i,0}$ to inherit its discriminative capability w.r.t. the nearest prototype $\boldsymbol{m}^{m}_{i,0}$.
Separation Loss: The basic separation loss $\mathcal{L}_{sep}$ is to minimize $l_{i,0}$ for normal samples while simultaneously maximizing them for anomalous instances, following the semi-supervised anomaly detection approach [29]:
@@ 3.5.1 Discrimination Loss
where $m_{0}$ prevents excessively large anomaly scores.
Margin Loss: The distribution overlap between normal and anomalous samples is also crucial for anomaly detection [6], which is considered by the margin loss $\mathcal{L}_{mar}$ as:
@@
where $m_{l}=\min_{i|y_{i}=1}l_{i,0}$ is the lower bound of $l_{i,0}$ for anomalous samples and $m_{u}=\max_{i|y_{i}=0}l_{i,0}$ is the upper bound for normal samples.
3.5.2 Geometry-Preserved Loss
Since learning a robust and generalizable $l(\cdot)$ in the high-dimensional space for anomaly detection is challenging, despite the architecture design in Section 3.4, we aim to exploit the desired geometric properties of $l(\cdot)$ that benefit our task.
Consistency Loss: Since Eqs. (4) and (5) focus on the discriminative power of $l_{i,0}$, we exploit the geometric consistency of $l_{i,j}$ across different local spaces around $\mathbf{m}^{m}_{i,j}$, including the following two properties:
@@ 3.5.2 Geometry-Preserved Loss
Accordingly, we formulate the consistency loss $\mathcal{L}_{cns}$ as:
@@
Despite the spatial regularization, $\mathcal{L}_{cns}$ also effectively increases sample diversity within each local space through multi-prototype observations as data augmentation.
@@
$\mathcal{L}_{sc}$ asymmetrically compresses scaling factors of normal samples while confining those of synthetic anomalies near the upper limit $e^{1}$. This design addresses the more pronounced deviations of synthetic anomalies than real-world ones, imposing more penalties on them.
Crossmodal Alignment Loss: Currently, both prototypes and directions are naively concatenated as input to LSPN (Sec. 3.4), ignoring their crossmodal correspondence. To resolve this, we propose a simple strategy by disentangling this property.
Specifically, by permuting indices $\{i\}\rightarrow\{\tilde{i}\}$, we can construct two kinds of negative samples:
@@
The former breaks holistic prototype coherence. In contrast, the latter focuses on capturing discordant directions while maintaining prototype consistency at a local structural granularity.
During training, we construct the above two kinds with equal probability. Denote $\tilde{w}^{m}_{i,0}$ ($m\in\{P,R\}$) as their associated scaling factors, we define the crossmodal alignment loss as:
@@ 3.5.3 Overall Loss
3.6 Geometry-Informed Anomaly Scoring
@@
This definition stems from a geometric prior. Specifically, memory bank-based methods calculate anomaly scores using Euclidean distances to their memory banks (Section 3.2), whereas our formulation in Eq. (11) reformulates this as distance measurement to the normal data manifold through the anisotropic metrics $l(\cdot)$. We will compare Eq. (11) with several alternatives in Section 4.3.
@@ 3.7 Anomaly Synthesis
DAUP [23] 0.976 0.977 0.980 0.960 0.924 0.966 0.981 0.978 0.972 0.980 0.969
Ours 0.982 0.977 0.982 0.979 0.971 0.976 0.982 0.983 0.978 0.981 0.979
Table 1: I-AUROC and AUPRO@30% results on MVTec-3D AD. Best results in bold, runner-ups underlined.
4 Experiment
@@
Eyecandies AST [28] 0.780 0.902 0.744 0.149
M3DM [32] 0.882 0.977 0.887 0.331
CFM [11] 0.881 0.973 0.887 0.335
2M3DF [1] 0.897 - 0.890 -
LSFA [30] - 0.974 - -
Ours 0.902 0.982 0.898 0.357
Table 2: Mean metrics for all categories on both datasets.
4.1 Experimental Details
Dataset: Experiments are conducted on the MVTec-3D AD [4] and Eyecandies[5] datasets. MVTec-3D AD contains 10 industrial categories (210-361 train, 69-132 test per class), while Eyecandies provides 10 synthetic food items (1k train, 400 test per class). Each sample provides aligned RGB images and 3D point clouds, enabling pixel-wise correspondence between color and spatial coordinates.
Evaluation Metrics: We assess the image-level anomaly detection performance by the Area Under the Receiver Operator Curve (I-AUROC). For evaluating the pixel-wise anomaly segmentation, we adopt the pixel-level AUROC (P-AUROC) and the Area Under the Per-Region Overlap (AUPRO) at 30% integration limits (denoted as AUPRO@30%), the same as existing methods [11, 19, 10, 32]. To better demonstrate the advantages of our method in practical industrial scenarios with lower false positive rates, we also adopted the AUPRO@1% metric (additional AUPRO@5% and AUPRO10% in the supplementary material) used in recent studies [10, 11].
Implementation: We follow a similar procedure of M3DM [32] for feature extraction and preprocessing of the dataset. For $\mathcal{L}_{cns}$, we set $k=5$ and $\eta_{0}=1.2$. We train our G2SF with the Adam optimizer [21] with 80 epochs on Linux Nvidia RTX 4090. More details are provided in the supplementary material.
Figure 6: Pixel-wise anomaly score distributions in Bagel, Cable gland, Foam, and Rope categories, where the horizontal axis is normalized score (0-1) with vertical axis as the probabilistic density. Green/Brown distributions represent normal/anomalous regions respectively. Compared with CFM [11] (top), our method (bottom) achieves significant distribution compression in normal regions while maintaining anomaly discrimination.